What is the equation of the line that has slope -4 and y-intercept (0,3)?
Found 3 solutions by jim_thompson5910, richard1234, richwmiller:
Answer by jim_thompson5910(35256)

y - y1 = m(x - x1)

y - 3 = -4(x - 0)

y - 3 = -4x

y = -4x + 3

Answer by richard1234(7193)

y = -4x + 3
Alternatively, you can write the point-slope form of the equation of the line.

Answer by richwmiller(17219)

y=mx+b
m=-4
b=3
y=-4x+3